# Bayesian sample size calculations for external validation studies of risk prediction models

### v1

@@ 3.0.2 Application of Bayesian approach
EVSI quantifies the expected value of a validation sample of a given size to identify the strategy with highest clinical utility
From this figure, one can (subjectively) realize that the CIW components fall on the ‘diminishing return’ of the EVSI. For example, going from the the largest sample size based on criteria other than the calibration slope (N=516) to the sample size dictated by assurance on calibration (N=1,173), a more than doubling of the sample size, is associated with 9.2% expected gain in NB. This can justify relaxing the criteria on calibration sloe for the validation study, if the focus is on clinical utility.
The decision whether to relax the calibration slope criteria can also be examined via the stability of the flexible calibration plots. As the desired precision based on the calibration slope is difficult to justify, but often dictates the final sample size required, Riley et al. recommend to plot the corresponding variability in calibration curves that might arise for that sample size33. In our context, the variability of a calibration curves has two sources: the variability in the true calibration function (due to variability in $\theta$) and the variability due to the finite validation sample (due to variability of $D_{N}$). The former is not a function of sample size. Because in our Bayesian Monte Carlo sampling algorithm, the true calibration function is known within each iteration, we can remove variability due to $\theta$ to quantify the difference between the flexible calibration curve and the true calibration function: for each simulated $D_{N}$, we fit a kernel regression (e.g., loess) for $P(Y=1|\pi)$, then remove $h(\pi)$ from the fitted values. Figure 4 shows such Bayesian calibration error plots for the final sample sizes including calibration slope criteria (N=1,173) and after removing such criteria (N=516). Moving to the smaller sample size expectedly increases the spread of calibration errors. Nevertheless, aside from extreme values of predicted risk, the 95% credible intervals (dashed lines) indicate that the error will mostly be <0.1. Again, this can be taken as a further argument in favor of settling on the smaller sample size.
@@ 4 Discussion
We proposed a Bayesian framework for sample size considerations when a risk prediction model is to be evaluated (validated) in a target population. Compared to the conventional frequentist framework, this approach offers several advantages: it allows investigators to specify their uncertainties about model performance; it enables assurance-type sample size rules that incorporate investigators’ risk preferences; and it facilitates the the use of Value-of-Information analysis when clinical utility of the model is being examined. We proposed characterizing uncertainties around common metrics of model performance, and introduced Monte Carlo sampling algorithms for computing precision and VoI metrics.
@@
Overall, this framework can provide an objective first step when designing validation studies aimed at quantifying the performance of a pre-specified model in a new population and determining whether it provides clinical utility. In this context, this framework helps the investigator explore the trade-off between sample size, precision, and gain in clinical utility. This framework can be used to investigate the anticipated precision or expected gain in clinical utility when the sample size is fixed (based on already collected data). Moroever, it can also be used to determine the sample size when original data collection is planned. A hybrid use, determining the sample size based on certain criteria and examining the resulting precision and VoI for other outcomes, is also possible. Our case study demonstrated such a hybrid approach. We defined rules targeting expected value and 90% assurance on the widths of confidence intervals around metrics of discrimination and calibration, as well as 90% assurance that we will be able to correctly identify the strategy with the highest clinical utility. Then, based on the results of the VoI analysis (as well as visual inspection of calibration errors), we could argue that one could potentially relax one component of the samples size requirement (precision around calibration slope), without losing much precision or clinical utility.
@@
Among the outputs of this Bayesian approach, the assurance-type rules and VoI metrics are particularly relevant for risk prediction modeling. If the estimates of model performance are too uncertain, the adoption of the model might be met with resistance, even if the point estimates are compatible acceptable performance. This, combined with the general risk-averse attitude of public funding agencies34, means the costs of not meeting the desired targets is typically more than the benefits of exceeding those targets. Assurance-type rules can be used alongside those based on expected values to assuage concerns that the planned study might generate ambiguous results. Further, VoI metrics address longstanding criticisms on the irrelevance of inferential statistics when deciding on whether to adopt a health technology (including markers and risk scoring tools) for patient care9. These criticism have been recently voiced in the particular case of NB calculations for risk prediction models10. Accordingly, we recommend sample size considerations around NB to move away from confidence intervals and focus on assurance and EVSI. In particular, EVSI combines the risk of failing to identify the optimal strategy with the consequences (NB losses) of such a failure, providing a fuller picture of the implications of learning from an empirical study. We note that EVSI captures the expected gain in NB per decision. When scaled to the population (i.e., multiplied by the expected number of time the decision of interest is to be made), it quantifies the total value of conducting a validation study of a given size in net true or false positive units (an example of such calculations are provided in a previous study18). This provides a ‘value-based’ perspective that can complement the precision-based approach for sample size determination around metrics of discrimination and calibration.
@@
There are several areas for further inquiry. We focused on binary outcomes, but this methodology can be extended to other outcome types. The application to survival outcomes seems particularly relevant and feasible. For such outcomes, model performance needs to be assessed at a time point of interest35. The time-dependent equivalents of prevalence and c-statistic are, respectively, the complement of survival probability and time-dependent area under the receiver operating characteristic curve at the time horizon of interest36. If uncertainty around these metrics are specified, our identifiability conditions can then be employed to map such metrics to the distribution of calibrated risks at this time point. However, the impact of censored observations in the future sample and the presence of competing risks on precision metrics and VoI values remains to be investigated. We focused on precision and VoI targets that pertain to the population-level performance of the model. Sample size consideration can also be approached in terms of uncertainty at the individual level, such as instability at individual-level predictions37. In line with Riley equations, we modeled the calibration function to be linear on the logit scale. This can be relaxed, for example by introducing a quadratic term as long as one can assure monotonicity of $h()$ and specify the joint distribution of the three parameters that would define it. A more appealing extension would be to model $h()$ non-parametrically. This can be done based a non-parametric summaries of calibration curves, such as those based on kernel smoothing (e.g., $ICI$16) or cumulative plots (e.g., $C^{*}$38).
@@
From an implementation perspective, Bayesian calculations are inherently more complex than their frequentist counterparts. Several components of this framework require numerical integration and optimization methods. Examples include finding the parameters of $P(p)$ given prevalence and c-statistic, mapping O/E ratio or calibration mean to calibration intercept, and computing sensitivity and specificity (for NB calculations) at the threshold of interest given $P(p)$ and $h()$. These numerical algorithms might struggle for extreme cases. In addition, as the Bayesian Monte Carlo sampling produces draws from the posterior distributions of precision targets (e.g., CI widths), sample size determination requires stochastic optimization techniques, for example finding the $N$ that corresponds to the first moment or a quantile of this distribution. These algorithms demand convergence assessment and if needed repetition with different starting values. While our implementation performed robustly in our examinations (including the case study), numerical accuracy and stability of these algorithms should be considered across a range of input values in dedicated studies. We consider the accompanying software to be an evolving implementation, subject to revisions and improvements.
@@
Compared to the conventional sample size determination methods, the Bayesian approach is predicated on one crucial extra step: characterizing our uncertainties about model performance. Such characterization is not currently a common practice in applied prediction modeling. The conventional wisdom in risk modeling studies is to report a ‘final model’ as a deterministic function that maps patient characteristics to an exact value of conditional outcome risk. This practice ignores the fact that predictions for a model trained using a finite sample are inherently uncertain. Differences across populations in the relationship between predictors and the outcome, differences in predictor assessment approaches across settings, and variations in outcome ascertainment methods further add to our uncertainties39. Fortunately, there have been recent calls on the importance of characterizing and communicating uncertainty around model predictions40. These calls encourage investigators to fully document and communicate uncertainty in model coefficients in reports of model development (e.g., reporting the covariance matrix of model coefficients, or reporting the coefficients of models fitted in bootstrapped copies of the original sample)40. These can further facilitate incorporating uncertainties into the design of subsequent studies.
@@
A thought-provoking consequence of adopting a Bayesian approach is the ultimate fate of the prior information. Reflecting current practices, we assumed the future validation sample will be the sole source of evidence on model performance once it is procured, in effect discarding our current knowledge once it is used to construct $P(\theta)$. But, if our existing knowledge is informative enough to warrant investigating the model in a new setting, should it not be used alongside the sample to inform our final judgment? Taking our case study as an example, the information on prevalence alone (based on the moments of its distribution) is equal to having learned its value from a sample of 280 individuals - a non-trivial amount of information! Incorporating such prior knowledge into what we learn from the data will reduce prediction uncertainty, which can affect both patient care (more precise predictions) and the design of empirical studies (requiring smaller samples to achieve the same targets). Of course, a perceived drawback is the subjectivity inherent in real-world Bayesian reasoning. Our case study might be seen as a stylized setup where predictive distributions from a meta-analysis were available and the assumption of exchangability of populations seemed tenable. Things can be more subjective, for example, when a single development study in another population is the sole source of evidence. Our current response to the fear of compromised objectivity is to completely drop the existing information. Perhaps one can instead define objectivity in terms of explicit and transparent specification of prior evidence and steps taken to account for population heterogeneity.
@@
5 Appendix
5.1 Appendix 1: Algorithm for parametric bootstrapping
1. For i in 1 to M (size of the Monte Carlo simulation) (a) Generate $D_{n}:=\{(\pi_{i},Y_{i})\}_{i=1}^{n}$, an iid sample of size $n^{\dagger}$ consisting of predicted risks and response values using the point estimates $\hat{\theta}=\{\hat{\phi},\hat{c},\hat{h}\}^{\ddagger}$. (b) Estimate $\theta^{(j)}=\{\phi^{(j)},c^{(j)},h^{(j)}\}$ from the above sample. 2. Estimate the correlation coefficients for $\theta$s. Generally, correlation coefficients are not sensitive to the size of the sample. It is natural to choose $n$ to be the size of the original sample. In instances (as in our case study) where elements of $\theta$ are taken from different sources or from the output of meta-analysis (as in our case study), one can determine $n$ based on the effective sample size (e.g., based on uncertainty around prevalence). DEtermining the distribution of predicted risks and observed responses from .
5.2 Appendix 2: Sample size and CI widths for the two-level closed-form method
Approach c-statistic O/E ratio calibration slope
Bayesian (expected CI width) 343 420 1044
Bayesian (90% assurance) 398 516 1173
Expected CI widths (a) 90th quantile of CI width (b), optimality assurance (c) curves based on the two-level closed-form method. Shapes on the line pertain to individual sample size components.
Refer to caption Refer to caption
(a) Expected value of CI width (b) 90th percentile of CI widths

### v2

@@ Methods
From this figure, one can (subjectively) realize that some of the CIW components fall on the ‘diminishing return’ of the EVSI curve. For example, going from the the largest sample size based on criteria other than the calibration slope (N=522) to the sample size dictated by assurance on calibration (N=1,181), a more than doubling of the sample size, is associated with only 8.8% expected gain in NB. This can justify relaxing the criteria on calibration slope for the validation study, if the focus is on clinical utility.
The decision whether to relax the calibration slope criteria can also be examined via the stability of the flexible calibration plots. As the desired precision based on the calibration slope is difficult to justify, but often dictates the final sample size required, Riley et al recommend plotting the corresponding variability in calibration curves that might arise for that sample size[9]. In our context, the variability of a calibration curves has two sources: the variability in the true calibration function (due to variability in $\theta$) and the variability due to the finite validation sample (due to variability of $D_{N}$). The former is not a function of sample size. Because in our Bayesian Monte Carlo sampling algorithm, the true calibration function is known within each iteration, we can remove variability due to $\theta$ to quantify the difference between the flexible calibration curve and the true calibration function: for each simulated $D_{N}$, we fit a kernel regression (e.g., LOESS) for $P(Y=1|\pi)$, then remove $h(\pi)$ from the fitted values. Figure 4 shows such Bayesian calibration error plots for the final sample sizes including calibration slope criteria (N=1,181) and after removing such criteria (N=522). Moving to the smaller sample size expectedly increases the spread of calibration errors. Nevertheless, aside from extreme values of predicted risk, the 95% credible intervals (dashed lines) indicate that the error will mostly be <0.1. Again, this can be taken as a further argument in favor of settling on the smaller sample size. Putting all these together, the final recommended sample size will be N=522.
@@ Discussion
We proposed a Bayesian framework for sample size considerations when a risk prediction model is to be evaluated (validated) in a target population. Compared to the conventional frequentist framework, this approach offers several advantages: it allows investigators to specify their uncertainties about model performance; in addition to sample size rules targeting expected CI widths, it enables assurance-type sample size rules that incorporate investigators’ risk preferences; and it facilitates the the use of Value of Information analysis when clinical utility of the model is being examined. We proposed characterizing uncertainties around common metrics of model performance, and introduced Monte Carlo sampling algorithms for computing precision and VoI metrics. We implemented this approach as the bayespmtools R package, with two functions, one for quantifying the anticipated precision / VoI values given a fixed sample size, and one for determining the sample size given pre-specified rules.
@@
Overall, this framework can provide an objective first step when designing validation studies aimed at quantifying the performance of a pre-specified model in terms of trade-off between sample size, precision, and gain in clinical utility. This framework can be used to investigate the anticipated precision or expected gain in clinical utility when the sample size is fixed (based on already collected data). It can also be used to determine the sample size when original data collection is planned. A hybrid use, determining the sample size based on certain criteria and examining the resulting precision and VoI for other outcomes, is also possible. Our case study demonstrated such a hybrid approach. We defined rules targeting expected value and 90% assurance on the widths of confidence intervals around metrics of discrimination and calibration, as well as 90% assurance that we will be able to correctly identify the strategy with the highest clinical utility. Then, based on the results of the VoI analysis (as well as visual inspection of calibration errors), we could argue that one could potentially relax one component of the samples size requirement (precision around calibration slope), without losing much precision or clinical utility.
@@
Among the outputs of this Bayesian approach, the assurance-type rules and VoI metrics are particularly relevant for risk prediction modeling. If the estimates of model performance are too uncertain, the adoption of the model might be met with resistance, even if the point estimates are compatible with acceptable performance. This, combined with the general risk-averse attitude of public funding agencies[37], means the costs of not meeting the desired targets is typically more than the benefits of exceeding those targets. Assurance-type rules can be used alongside those based on expected values to assuage concerns that the planned study might generate ambiguous results. Further, VoI metrics address longstanding criticisms on the irrelevance of inferential statistics when deciding on whether to adopt a health technology (including markers and risk scoring tools) for patient care[10]. These criticism have been recently voiced in the particular case of NB calculations for risk prediction models [11]. Accordingly, we recommend sample size considerations around NB to move away from confidence intervals and focus on assurance and EVSI. In particular, EVSI combines the risk of failing to identify the optimal strategy with the consequences (NB losses) of such a failure, providing a fuller picture of the implications of learning from an empirical study in terms of clinical utility. This provides a ‘value-based’ perspective that can complement the precision-based approach for sample size determination around metrics of discrimination and calibration.
@@
There are several areas for further inquiry. We focused on binary outcomes, but this methodology can be extended to other outcome types. The application to survival outcomes seems particularly relevant and feasible. For such outcomes, model performance needs to be assessed at a time point of interest[38]. The time-dependent equivalents of prevalence and c-statistic are, respectively, the complement of survival probability and time-dependent area under the Receiver Operating Characteristic curve[39]. If uncertainty around these metrics are specified, our identifiability conditions can then be employed to map such metrics to the distribution of calibrated risks at this time point. However, the impact of censored observations in the future sample and the presence of competing risks on precision metrics and VoI values remains to be investigated. We focused on precision and VoI targets that pertain to the population-level performance of the model. Sample size consideration can also be approached in terms of uncertainty at the individual level, such as instability at individual-level predictions[40]. Further, when contemplating a validation study, often a secondary objective is to update (revise) the model if its performance turns out to be sub-optimal. It makes sense to consider targets related to model updating when designing validation studies[41]. In line with Riley equations, we modeled the calibration function to be linear on the logit scale. This can be relaxed, for example by introducing a quadratic term as long as one can assure monotonicity of $h()$ and specify the joint distribution of the three parameters that would define it. A more appealing extension would be to model $h()$ non-parametrically. This can be done based a non-parametric summaries of calibration curves, such as those based on kernel smoothing (e.g., $ICI$[17]) or cumulative plots (e.g., $C^{*}$[42]).
@@
From an implementation perspective, Bayesian calculations are inherently more complex than their frequentist counterparts. Several components of this framework require numerical integration and optimization methods. Examples include finding the parameters of $P(p)$ given prevalence and c-statistic, mapping O/E ratio or calibration mean to calibration intercept, and computing sensitivity and specificity (for NB calculations) at the threshold of interest given $P(p)$ and $h()$. These numerical algorithms might struggle for extreme cases. In addition, as the Bayesian Monte Carlo sampling produces draws from the pre-posterior distributions of precision targets (e.g., CI widths), sample size determination requires stochastic optimization techniques. These algorithms demand convergence assessment and if needed repetition with different starting values. While our implementation performed robustly in our examinations (including the case study), numerical accuracy and stability of these algorithms should be considered across a range of input values in dedicated studies. We consider the accompanying software to be an evolving implementation, subject to revisions and improvements.
@@
Compared to the conventional sample size determination methods, the Bayesian approach is predicated on one crucial extra step: characterizing our uncertainties about model performance. Such characterization is not currently a common practice in applied prediction modeling. The conventional wisdom in risk modeling studies is to report a ‘final model’ as a deterministic function that maps patient characteristics to an exact value of conditional outcome risk. This practice ignores the fact that predictions for a model trained using a finite sample are inherently uncertain. Differences across populations in the relationship between predictors and the outcome, differences in predictor assessment approaches across settings, and variations in outcome ascertainment methods further add to our uncertainties[43]. Fortunately, there have been recent calls on the importance of characterizing and communicating uncertainty around model predictions[44]. These calls encourage investigators to fully document and communicate uncertainty in model coefficients in reports of model development (e.g., reporting the covariance matrix of model coefficients, or reporting the coefficients of models fitted in bootstrapped copies of the original sample)[44]. These can further facilitate incorporating uncertainties into the design of subsequent studies.
@@
A thought-provoking consequence of adopting a Bayesian approach is the ultimate fate of the prior information. Reflecting current practices, we assumed the future validation sample will be the sole source of evidence on model performance once it is procured, in effect discarding our current knowledge once it is used to construct $P(\theta)$. But, if our existing knowledge is informative enough to warrant investigating the model in a new setting, should it not be used alongside the sample to inform our final judgment? Taking our case study as an example, the information on prevalence alone (based on the moments of its distribution) is equal to having learned its value from a sample of 281 individuals - a non-trivial amount of information! Incorporating such prior knowledge into what we learn from the data will reduce prediction uncertainty, which can affect both patient care (more precise predictions) and the design of empirical studies (requiring smaller samples to achieve the same targets). Of course, a perceived drawback is the subjectivity inherent in real-world Bayesian reasoning. Our case study might be seen as a stylized setup where predictive distributions from a meta-analysis were available and the assumption of exchangability of populations seemed tenable. Things can be more subjective, for example, when a single development study in another population is the sole source of evidence. Our current response to the fear of compromised objectivity is to completely drop the existing information. Perhaps one can instead define objectivity in terms of explicit and transparent specification of prior evidence and steps taken to account for population heterogeneity.